tanA is the integral solution of 
inequality 4x^2-16x+15
cosB is the slope of bisector
of the first quadrant then the 
value of sin(A+B)*sin(A-B) is

Grade:12

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

Solution of 4x^2-16x+15=0
factorize we get (2x-3)(2x – 5)=0
So roots are x= 5/2 and 3/2 => tanA ….…..(1)
If you know tanA you can find out all other trigonometric functions

cosB= slope of bisectorof the first quadrant = tan45 = 1
B= 0 degrees
sin(A+B)*sin(A-B)= sin 2A [as B=0]
Now use equation (1) and find the answer